Let’s start with the simplest valuation model available to us: the perpetuity formula. Suppose you expect...

  1. Let’s start with the simplest valuation model available to us: the perpetuity formula.

Suppose you expect a company to generate FCF to the Firm of $100m next year, expect it to grow at a sustainable growth rate of 2.5%, and believe that it has a Weighted Average Cost of Capital (WACC) of 5%.

You also know that this firm has $1,500m in debt outstanding, and that there are 100 shares of its stock outstanding.

  1. What is the value of this firm using the assumptions provided above, and what is the value of equity implied by your estimate?
Answer #1

Expected FCF = FCF1 = $ 100 million, Growth Rate = g = 2.5 % and Weighted Average Cost of Capital (WACC) = 5 %

Debt = $ 1500 million and Number of Shares Outstanding = N = 100 mllion

Firm Value = FCF1 / (WACC - g) = 100 / (0.05 - 0.025) = $ 4000 million

Equity Value = 4000 - 1500 = $ 2500 million

Share Price = 2500 / 100 = $ 25